OpenStudy (anonymous):

Could some one help me with every step to this algebra equation b-9+6b=30

OpenStudy (cwrw238):

aim to get terms in b to one side of equal sign and numbers to the right to get -9 to the right add 9 to both sides b - 9 + 9+ 6b = 30 + 9 b + 6b = 39 ( 9's disappear on left hand side) 7b = 39 we need to find the value of b so we divide both sides b 7 7b/7 = 39/7 b = 39/7 or 5 4/7

OpenStudy (anonymous):

Okay I got thay far but I don't know what to do to check it as in plug it in .